# Anderson localization for the multi-frequency quasi-periodic CMV matrices and quantum walks

### v1

@@ 5.1 Gesztesy-Zinchenko cocycle
$C_{n}=C_{n,x,\omega}=\left[\begin{array}[]{cc}c_{n}^{11}(x+n\omega)&c_{n}^{12}(x+n\omega)\\ c_{n}^{21}(x+n\omega)&c_{n}^{22}(x+n\omega)\end{array}\right]$
with analytic sampling functions and $\omega\notin\mathbb{Q}$. We denote the corresponding quantum walk by $U_{\omega}(x)$.
As for the extended CMV matrix $\hat{\mathcal{E}}$, there is a decomposition $\hat{\mathcal{E}}=\hat{\mathcal{L}}\hat{\mathcal{M}}$ (see section 3.4 for details). For a given solution $u$ of ^$\hat{\mathcal{E}}u=zu$, we define $v=\hat{\mathcal{L}}^{-1}u$. One can verify that ^$\hat{\mathcal{M}}u=zv$. Then according to [13, section 2], we have that
$\left[\begin{array}[]{c}u(n+1)\\ v(n+1)\end{array}\right]=Y(n,z)\left[\begin{array}[]{c}u(n)\\ v(n)\end{array}\right],$
@@
$Y(n,z)=\frac{1}{\rho_{n}}\begin{cases}\left[\begin{array}[]{cc}-\hat{\alpha}_{n}&1\\ 1&-\overline{\hat{\alpha}}_{n}\end{array}\right]\quad n\,\,\text{is even},\\ \\ \left[\begin{array}[]{cc}-\overline{\hat{\alpha}}_{n}&z\\ z^{-1}&-\hat{\alpha}_{n}\end{array}\right]\quad n\,\,\text{is odd}.\end{cases}\$
For the sake of convenience, we write $Y(n,z)=P(\hat{\alpha}_{n},z)$ when $n$ is even and $Y(n,z)=Q(\hat{\alpha}_{n},z)$ when $n$ is odd. According to [8, section 3], we can define Gesztesy-Zinchenko cocycle by
@@
$G(\omega,z;x)=\frac{1}{\rho(x)}\left[\begin{array}[]{cc}z&-\overline{\hat{\alpha}}(x)\\ -\hat{\alpha}(x)&z^{-1}\end{array}\right].$
@@
5.2 Anderson localization for the quantum walks
According to the analysis in section 2, the Szegő cocycle and the $n$-step transfer matrix of the matrix $\hat{\mathcal{E}}$ are
$\hat{M}(\omega,z;x)=\frac{1}{\rho(x)}\left[\begin{array}[]{cc}\sqrt{z}&-\frac{\overline{\hat{\alpha}}(x)}{\sqrt{z}}\\ -\hat{\alpha}(x)\sqrt{z}&\frac{1}{\sqrt{z}}\end{array}\right]$
@@ 5.2 Anderson localization for the quantum walks
$G(\omega,z;x)=\left[\begin{array}[]{cc}\frac{1}{\sqrt{z}}&0\\ 0&\sqrt{z}\end{array}\right]\hat{M}(\omega,z;x).$
@@
$\left[\begin{array}[]{cc}\frac{1}{\sqrt{z}}&0\\ 0&\sqrt{z}\end{array}\right]=:M_{0}.$
@@
[…]
$=\lim_{n\rightarrow\infty}\frac{1}{n}\int_{\mathbb{T}^{d}}\log(C_{\mathrm{nor}})^{n}+\log\|\hat{M}_{n}(\omega,z;x)\|)dx$
$=\log C_{\mathrm{nor}}+\hat{L}(w,z).$
On the other hand, when $z=1$, $G_{n}(\omega,z;x)=\hat{M}_{n}(\omega,z;x)$. It follows that
$L^{U}(w,z)\geq\hat{L}(w,z).$
From the above discussion, quantum walks can be viewed as a special class of extended CMV matrices whose Veblunsky coefficients with even index are equal to zero. Due to the relationship between the quantum walks and the extended CMV matrices, we can get the Anderson localization for the quantum walks with multi-frequency quasi-periodic coins.
Theorem 5.1.
The family of multi-frequency quasi-periodic quantum walks $\{U_{\omega}(x)\}$ with analytic coins display Anderson localization in the regime of positive Lyapunov exponents for a fixed phase vector $x_{0}$ and almost all frequency vectors satisfying the standard Diophantine condition.
Remark 5.2.
This statement is an application of Theorem 1.1. Since
$\hat{L}(w,z)\leq L^{U}(w,z)\leq\log C_{\mathrm{nor}}+\hat{L}(w,z),$
then the Anderson localization result for the quantum walks follows.

### v3

@@ 4.2 Gesztesy-Zinchenko Cocycle
$C_{n}=C_{n,x,\omega}=\left(\begin{array}[]{cc}c_{n}^{11}(x+n\omega)&c_{n}^{12}(x+n\omega)\\ c_{n}^{21}(x+n\omega)&c_{n}^{22}(x+n\omega)\end{array}\right)$
with analytic sampling functions and frequency vector $\omega\in\mathbb{T}^{d}(p,q)$. Here we also assume that $\mathrm{det}\,C_{n,x,\omega}=1$. We denote the corresponding quantum walk by $U_{\omega}(x)$.
As for the extended CMV matrix $\hat{\mathcal{E}}$, there is a decomposition $\hat{\mathcal{E}}$$\hat{\mathcal{L}}\hat{\mathcal{M}}$ (see Section 2.4 for details). For a given solution $u$ of $\hat{\mathcal{E}}u=zu$, we define $v=\hat{\mathcal{L}}^{-1}u$, then $\hat{\mathcal{M}}u=zv$. According to [17, Section 2], we have that
$\left(\begin{array}[]{c}u(n+1)\\ v(n+1)\end{array}\right)=Y(n,z)\left(\begin{array}[]{c}u(n)\\ v(n)\end{array}\right),$
@@
$Y(n,z)=\frac{1}{\rho_{n}}\begin{cases}\left(\begin{array}[]{cc}-\hat{\alpha}_{n}&1\\ 1&-\overline{\hat{\alpha}}_{n}\end{array}\right)\quad n\,\,\text{is even},\\ \\ \left(\begin{array}[]{cc}-\overline{\hat{\alpha}}_{n}&z\\ z^{-1}&-\hat{\alpha}_{n}\end{array}\right)\quad n\,\,\text{is odd}.\end{cases}\\$
For the sake of convenience, we write $Y(n,z)=P(\hat{\alpha}_{n},z)$ when $n$ is even and $Y(n,z)=Q(\hat{\alpha}_{n},z)$ when $n$ is odd. According to [12, Section 3], we can define Gesztesy-Zinchenko cocycle by
@@
$G(\omega,z;x)=\frac{1}{\rho(x)}\left(\begin{array}[]{cc}z&-\overline{\hat{\alpha}}(x)\\ -\hat{\alpha}(x)&z^{-1}\end{array}\right).$
@@
To define the Lyapunov exponent of quantum walks, one can also consider other transfer matrices instead of Gesztesy-Zinchenko cocycle; for more details, see [9, 10]. Indeed, all these cocycles are equivalent.
Remark 4.1.
For split-step quantum walks, the transfer matrix in [10, Eq.(2.4)] provides a non-alternating representation, but our results hold under any equivalent cocycle.
4.3 Anderson Localization for Analytic Multi-Frequency Quantum Walks
According to the analysis in Section 1.3, the Szegő cocycle and the $n$-step transfer matrix of the matrix $\hat{\mathcal{E}}$ are
$\hat{M}(\omega,z;x)=\frac{1}{\rho(x)}\left(\begin{array}[]{cc}\sqrt{z}&-\frac{\overline{\hat{\alpha}}(x)}{\sqrt{z}}\\ -\hat{\alpha}(x)\sqrt{z}&\frac{1}{\sqrt{z}}\end{array}\right)$
@@ 4.3 Anderson Localization for Analytic Multi-Frequency Quantum Walks
$G(\omega,z;x)=\left(\begin{array}[]{cc}\frac{1}{\sqrt{z}}&0\\ 0&\sqrt{z}\end{array}\right)\hat{M}(\omega,z;x).$
@@
$\left(\begin{array}[]{cc}\frac{1}{\sqrt{z}}&0\\ 0&\sqrt{z}\end{array}\right)=:M_{0}.$
@@
$\leq\lim_{n\rightarrow\infty}\frac{1}{n}\int_{\mathbb{T}^{d}}\log((C_{\mathrm{nor}})^{n}\cdot\|\hat{M}_{n}(\omega,z;x)\|)dx$
$=\log C_{\mathrm{nor}}+\hat{L}(\omega,z).$
Since $|z|=1$, then $\|M_{0}\|=1$ and $\log C_{\mathrm{nor}}=0$. Therefore, we have that
$L^{U}(\omega,z)\leq\hat{L}(\omega,z).$ (4.6)
To this end, building upon Theorem 1.1 and the unitary equivalence established in Section 4.1, we now can present the main localization result for quantum walks with multi-frequency quasi-periodic coins.
Theorem 4.2 (Anderson Localization for Quantum Walks).
Consider a quantum walk $U_{\omega}(x)$ on $\ell^{2}(\mathbb{Z})\otimes\mathbb{C}^{2}$ with coins $C_{n}(x)=C(x+n\omega),\omega\in\mathbb{T}^{d}(p,q)$ satisfying:
  1. 1.
    $C:\mathbb{T}^{d}\to U(2)$ is real-analytic with analytic extension to $\mathbb{T}_{h}^{d}$;
  2. 2.
    $\det C(x)=1$ for all $x\in\mathbb{T}^{d}$;
  3. 3.
    The Lyapunov exponent $L^{U}(\omega,z)>\gamma>0$ for all $\omega\in\mathbb{T}^{d}(p,q)$ and all $z\in\mathcal{I}_{C}\subset\partial\mathbb{D}$, $\mathcal{I}_{C}$ is a compact interval.
For fixed $x_{0}\in\mathbb{T}^{d}$ and almost every $\omega\in\mathbb{T}^{d}(p,q)$, the quantum walk exhibits AL. Specifically, there exists a complete set of eigenfunctions $\{\psi_{n}\}$ such that
$U_{\omega}(x_{0})\psi_{n}=e^{i\theta_{n}}\psi_{n}$
with $\theta_{n}\in[0,2\pi)$ being the eigenphases, and the eigenfunctions decay exponentially:
$|\psi_{n}(k)|\leq c_{n}e^{-\gamma|k|/4}\quad\mathrm{for}\,\,\mathrm{all}\,\,k\in\mathbb{Z},$
where $c_{n}$’s are some constants.
Moreover, the unitary operator $U_{\omega}(x_{0})$ has pure point spectrum in $\mathcal{I}_{C}\subset\partial\mathbb{D}$.
Proof.
The proof proceeds via the unitary equivalence established in Section 4.1.
From (4.4), we see that $U_{\omega}(x)$ is unitarily equivalent to the extended CMV matrix $\hat{\mathcal{E}}_{\omega}(x)$ via $\mathcal{D}$.
The Verblunsky coefficients $\hat{\alpha}_{n}(x)$ of $\hat{\mathcal{E}}_{\omega}(x)$ satisfy:
$\hat{\alpha}_{2n}=0,\quad\hat{\alpha}_{2n+1}=-\frac{\lambda_{2n+1}}{\lambda_{2n+2}}\cdot c_{n}^{12}(x)$
with $|\lambda_{j}|=1$, hence $\hat{\alpha}_{n}(x)$ are real-analytic on $\mathbb{T}^{d}$.
The Lyapunov exponent relation (4.6) gives that
$\hat{L}(\omega,z)\geq L^{U}(\omega,z)>\gamma>0.$
Then applying Theorem 1.1 to $\hat{\mathcal{E}}_{\omega}(x)$ we can conclude that: for a.e. $\omega\in\mathbb{T}^{d}(p,q)$, $\hat{\mathcal{E}}_{\omega}(x_{0})$ has AL in $\mathcal{I}_{C}$.
By unitary equivalence, the eigenfunctions $\phi_{n}$ of $\hat{\mathcal{E}}_{\omega}(x_{0})$ correspond to eigenfunctions $\psi_{n}=\mathcal{D}\phi_{n}$ of $U_{\omega}(x_{0})$ with identical decay rates.
The pure point spectrum follows from the completeness of eigenfunctions and unitary equivalence. ∎
@@
We extend our gratitude to Prof. Christopher Cedzich (Heinrich Heine Universität Düsseldorf) for his insightful suggestions on the comprehension of quantum walks, which greatly enhanced the rigorousness of our results.